OpenStudy (anonymous):

Subtracting Functions help if f(x)=3x-2 and g(x)=x^2-4x+1,find (f-g)(x)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I know the answer But I don't know how to get there

OpenStudy (anonymous):

First i would write equation in one line so that i can see what i am subtract 3x-2- (x^2 - 4x+1) Then i would distribute the negative sign so it would be 3x -2 -x^2 + 4x -1

OpenStudy (anonymous):

solve it that way.

OpenStudy (anonymous):

i guess then you can put the like terms together. so -x^2 +4x +3x -1 -2

OpenStudy (anonymous):

-x^2 + 7x -3 should be the answer. did you get that?

OpenStudy (anonymous):

I don't understand where -x^2 came from. Wouldn't 3x-x^2 be 2x^2?

OpenStudy (anonymous):

I think I need a equation calculator

OpenStudy (anonymous):

@bobristic

OpenStudy (anonymous):

no 3x-x^2 would not be 2x^2 because 3x is a term to the power of one (3x^1) and to subtract a variable you need to have the same powerand -x^2 has the power of two. the -x^2 came from when you distributed the - (x^2 - 4x+1). the negative outside of the parenthesis is distributed to all the terms in the parenthesis. x^2 was positive and because the - was distributed it was now x^2.

OpenStudy (anonymous):

what do you mean by equation calculator?

OpenStudy (anonymous):

Thats a subtraction sign not a negative sign

OpenStudy (anonymous):

you called it a negative sign...I think you maybe be confusing me or your confused

OpenStudy (anonymous):

well i know when you say 3x-2- (x^2 - 4x+1) it is just shorten version of writing 3x-2 + (-1)(x^2 - 4x+1)

OpenStudy (anonymous):

is it though in this case?Because I'm specifically subtracting

OpenStudy (anonymous):

what do you mean though? to be honest negative sign or subration is practically the same thing . Could you rewrite what you are confused about because im confused about what you are confused about.

OpenStudy (anonymous):

Ok hold on

OpenStudy (anonymous):

I just dont understand where -x^2 came from and "7x" came from when I'm supposed to be subtracting if f(x)=3x-2 and g(x)=x^2-4x+1 find (f-g)(x)

OpenStudy (anonymous):

okay so when you are finding (f-g)(x), you are subtracting f(x) - g(x). then you replace the function with the equation. 3x-2 - (x^2-4+1) do you understand so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay. then you got to make the parenthesis go away. but there is a subtraction/negative sign in front of it, there for you have to give that negative sign to every thing(terms) a negative symbol. 3x-2 -(x^2) -(-4x) -(1). and -(-) is just positive so 3x-2 -x^2 +4x -1. so far with me?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then you have to add all the like terms, which are number with variables with same same power. 3x has the power of 1. -2 has the power of 0 -x^2 has the power of 2 +4x has the power of 1 -1 has the power of 0

OpenStudy (anonymous):

then put the number in highest power to the lowest power -x^2 +3x+4x -2 -1

OpenStudy (anonymous):

then you can see the like terms together and you add/subtract the terms. so 3x+ 4x = 7x and -2-1= -3 then write down the equation with the term added. -x^2 +7x-3

OpenStudy (anonymous):

is there a simpler way to do this?

OpenStudy (anonymous):

no offense

OpenStudy (anonymous):

well this is simple ... if i was there physically i would be able to explain better..

OpenStudy (anonymous):